OpenStudy (anonymous):

medal for answer Choose the correct simplification of a to the 5th power times b to the 6th power all over a to the 4th power times b to the 4th power.

OpenStudy (anonymous):

a9b10 ab2 1 over a times b to the 2nd power 1 over a to the 9th power times b to the 10th power

OpenStudy (kirbykirby):

\[ \frac{a^5b^6}{a^4b^4}\]?

OpenStudy (anonymous):

a9b10

OpenStudy (kirbykirby):

\[ \frac{a^5b^6}{a^4b^4}=\frac{a^{5}}{a^4}\cdot\frac{b^6}{b^4}=a^{5-4}b^{6-4}\]

OpenStudy (anonymous):

a1b2

OpenStudy (kirbykirby):

yep

OpenStudy (anonymous):

so ts c

OpenStudy (anonymous):

can you help me with one more

OpenStudy (kirbykirby):

no. \(a^1=a\)

OpenStudy (anonymous):

so its c o__o ...

OpenStudy (kirbykirby):

\(a^1\ne\dfrac{1}{a}\)

OpenStudy (anonymous):

O NVM its B

OpenStudy (kirbykirby):

yes! hehe

OpenStudy (anonymous):

thanks and uhm,can you help me with this:Choose the correct simplification of x to the 2nd power over y to the 5th power all raised to the 4th power. x to the 8th power over y to the 9th power y12 x12y8 x to the 8th power over y to the 20th power i think its d

OpenStudy (kirbykirby):

yes indeed :)

OpenStudy (anonymous):

thanks

OpenStudy (kirbykirby):

your welcome